### v3

# SoliDiffy: AST Differencing for Solidity Smart Contracts

Abstract
Smart contracts, primarily written in Solidity, are integral to blockchain software applications, yet precise analysis and maintenance are hindered by the limitations of existing differencing tools. We introduce SoliDiffy, a novel Abstract Syntax Tree (AST) differencing tool specifically designed for Solidity. SoliDiffy enables fine-grained analysis by generating accurate and concise edit scripts of smart contracts , making it ideal for downstream tasks such as vulnerability detection, automated code repair, and code reviews. Our comprehensive evaluation on a large dataset of real-world Solidity contracts demonstrates that SoliDiffy delivers shorter and more precise edit scripts compared to state-of-the-art tools, while performing consistently in complex contract modifications. SoliDiffy is made publicly available at https://github.com/mojtaba-eshghie/SoliDiffy.
Index Terms:
AST Differencing, Solidity, Smart Contracts, Abstract Syntax Tree, Ethereum, Source Code Differencing
I Introduction
Smart contracts are self-executing programs that implement real-world contracts by encoding contract terms directly into code [1, 2]. These programs are deployed on blockchain platforms like Ethereum [3], allowing for automated and trustless transactions. Solidity, a statically-typed programming language, has become the most popular choice for developing these smart contracts [4].
Developers working with Solidity can greatly benefit from a fine-grained source code differencing in several scenarios. For instance, when updating a smart contract to patch a security vulnerability, Abstract Syntax Tree (AST)-based differencing allows developers to pinpoint specific changes in the code’s structure rather than sifting through line-by-line text changes, which might miss subtle yet crucial modifications. This precision is vital when reviewing updates for correctness and security implications before deployment on an immutable blockchain. Other examples are in automated debugging [5] and program repair [6, 7, 8, 9], where tools need to detect bugs and vulnerabilities [10, 11, 12, 13, 14] and suggest changes at a syntactic level [15] ; fine-grained differencing enables these tools to generate precise edit scripts that align with the semantic intentions of the code, helping with automated and context-aware fixes. Furthermore, in code clone detection [16, 17, 18], where identifying syntactically similar but not identical code blocks is necessary, AST differencing can accurately capture variations that line-based tools overlook, thus enhancing the detection of potential code reuse or duplication issues.
Traditional text and structural differencing tools [19, 20, 21, 22, 23, 24, 25] are insufficient for smart contracts due to their inability to capture the semantic and structural details of smart contracts or their lack of support for Solidity. AST differencing for Solidity smart contracts provides a more granular approach at detection of changes at the syntactic level. To summarize, we problem we address in this paper is the fundamental limitations of existing differencing tools for Solidity developers.
In this paper, we introduce SoliDiffy, a novel AST differencing tool tailored for Solidity smart contracts. SoliDiffy contains key AST transformations specifically designed for Solidity smart contracts. SoliDiffy accurately processes and compares Solidity smart contracts, providing precise edit scripts that can be used for tasks such as vulnerability detection.
To evaluate the effectiveness of SoliDiffy, we conduct a comprehensive comparison on 354 187354187354\,187354 187 pairs of smart contracts. This dataset is founded on real-world Solidity smart contracts, including a subset of modified controlled syntactic changes (simple to complex transformation) and another subset mined from commit history of a popular smart contract repository. We assess the performance of SoliDiffy in terms of edit script accuracy. The results clearly demonstrate that SoliDiffy outperforms the only existing tool for Solidity differencing, Difftastic [21]. SoliDiffy maintains consistent effectiveness regardless of edit distance and code complexity, highlighting its suitability for advanced software engineering tasks in the blockchain domain.
To summarize, our contributions are:
  • We introduce SoliDiffy, a novel AST differencing approach specifically designed for Solidity smart contracts, addressing the limitations of existing differencing tools in accurately capturing the syntactic and semantic changes within smart contracts.
  • We design Solidity-specific AST transformation and pruning rules that enhance the precision and conciseness of edit scripts compared to a raw concrete syntax tree.
  • We conduct a comprehensive evaluation of SoliDiffy on a dataset of 354 187354187354\,187354 187 pairs of smart contracts to diff, demonstrating that it outperforms the state-of-the-art tool Difftastic in terms of edit script length and accuracy. SoliDiffy maintains consistent effectiveness regardless of edit distance and code complexity.
  • We make SoliDiffy publicly available as an open-source tool at https://github.com/mojtaba-eshghie/SoliDiffy, facilitating further research and applications in smart contract analysis.
This paper is structured as follows: Section II introduces the necessary background concepts, including ASTs and code differencing techniques. In Section III, we describe the architecture of SoliDiffy. Section IV outlines the experimental protocol including research questions that guide our investigation. Section V presents the results of our evaluation, comparing SoliDiffy with Difftastic and exploring the impact of different types and severity of syntactic changes on the performance of these tools. Section VII reviews existing work related to AST differencing and identifies gaps that SoliDiffy addresses. Finally, Section VI elaborates the lessons learnt and threats to validity of our work, and Section VIII concludes the paper.
Figure 1: The design of the SoliDiffy smart contract differencing tool.
II Background
This section provides the concepts essential for AST-based code differencing.
II-A Syntax Trees: Abstract vs. Concrete
Syntax trees, comprising abstract and concrete syntax trees (ASTs and CSTs), represent the hierarchical structure of source code. ASTs focus on the logical structure by abstracting away syntactic details, making them ideal for tasks like code analysis and transformation [26]. In contrast, CSTs retain all syntactic elements, including punctuation and keywords, capturing the exact format of the source code, which is essential for precise replication tasks like formatting and refactoring.
II-B Code Differencing
Code differencing is the process of identifying differences between two versions of a codebase. This is crucial for version control, collaborative development, and maintaining code quality. Traditional line-based differencing tools, such as those used in Git [27], compare code on a line-by-line basis, which can miss or misinterpret finer structural changes in the code.
II-B1 AST Differencing
AST differencing enhances code comparison by utilizing the hierarchical structure of ASTs. Unlike line-based differencing, AST differencing can identify specific modifications within the code’s logical structure. For example, a small change within a line of code can be pinpointed precisely, rather than being treated as a completely new or altered line [19].
II-B2 Edit Scripts
A common approach in AST differencing involves generating an edit script as a sequence of operations required to transform one AST into another (see Figure 3). The process typically involves two phases: generating mappings between unchanged nodes of the two ASTs and then deriving an edit script from these mappings. These edit actions—add, remove, update, and move—reflect modifications to source code. Although generating an optimal edit script is an NP-hard problem [28], this method provides a structured way to represent differences between code versions.
III SoliDiffy : AST Differencing for Solidity
This section outlines the core components of SoliDiffy , a novel approach for fine-grained and precise AST differencing of Solidity smart contracts.
Figure 1 shows SoliDiffy’s architecture. SoliDiffy starts by receiving Solidity smart contracts and generating a optimized ASTs for the differencing task (Section III-A). The differencing subsystem then uses the mapping between the ASTs to perform the differencing (Section III-B).
III-A Pruning Rules
CSTs contain a wide range of unnecessary information that may pollute edit scripts. To create ASTs optimized for differencing, SoliDiffy  employs a series of transformations on the initial CSTs. This involves flattening nodes, aliasing for consistency, and pruning unnecessary elements, as follows
  • Flattening: Combines child nodes with their parent as a single node. In other words, we stop at one node in the AST and putting as value all the source string corresponding to this node and its children. As an example, the constant literal values with type and value are concatenated (first rule in under mapping rules and green sub-trees in Figure 1).
  • Aliasing: Renames node types to facilitate a unified differencing process (second rule “aliased” in Figure 1).
  • Ignoring: Removes extraneous nodes, such as formatting elements, that do not impact the logical structure of the code (third rule in Figure 1).
These transformation rules capture Solidity-specific constructs, and are implemented on top of pre-implemented transformation rules in Gumtree [19].
III-B Differencing Algorithm
The differencing algorithm of SoliDiffy  generates mappings between nodes of the two ASTs to identify unchanged and modified elements. Utilizing the efficient algorithm of Gumtree [19], which has undergone extensive evaluation, SoliDiffy  aligns nodes between two given ASTs. This process involves a two-phase mapping strategy:
  • Top-Down Mapping: Identifies large, unmodified subtrees to serve as anchors, reducing the complexity of subsequent differencing.
  • Bottom-Up Mapping: Refines the initial mappings by comparing smaller subtrees and individual nodes, ensuring that all modifications are accurately captured.
These mappings are then used to derive an edit script (Section III-C).
III-C Edit Script
SoliDiffy’s edit scripts include four standard operations: add, delete, update, and move. The differencing algorithm prioritizes producing edit scripts that are concise yet fully descriptive of the changes made.
To demonstrate edit scripts SoliDiffy generates, we use the smart contract in Figure 2. We change the identifier for a storage variable and remove a function from the source code. SoliDiffy generates the edit script in Figure 3 by identifying one update and a delete action.
1contract SimpleStorage {
2   - uint256 public num;
3   + uint256 public number_;
4 function set(uint256 _num) public {
5 num = _num;
6 }
7   - function get() public view returns (uint256) {
8   - return num;
9   - }
10}
Figure 2: Standard line diff of original and modified SimpleStorage contract, with added and removed lines highlighted.
Figure 3: Edit script generated as a result of the AST-based differencing of the code from Figure 2 using SoliDiffy. The edit actions are color-coded with delete actions presented as with red and update with orange.
III-D Implementation
SoliDiffy  uses Tree-sitter to parse Solidity source code into its syntax tree representation. Tree-sitter is an open-source parsing tool that generates CSTs for a wide range of programming languages using a modular grammar framework. We integrated the most recent version of Solidity grammar to the best of our knowledge [29]  [^1] https://github.com/JoranHonig/tree-sitter-solidity/blob/a8ed2f5d600fed77f8ed3084d1479998c649bca1/grammar.js.
IV Experimental Protocol
Figure 4: Pipeline for large-scale generation of contract pair for differencing and subsequent analysis.
We outline the experimental setup to answer our research questions 1–4 in Section IV-A. Figure 4 demonstrates this experimental setup, and Section IV-B elaborates on the protocol to answer the research questions.
The experiment is structured as a pipeline that begins with the selection and preparation of our seed datasets of Solidity smart contracts. The seed datasets include 1) DAppSCAN, a large dataset of smart contracts [30] and seed dataset from commit history of the Uniswap smart contracts [31] (Section IV-C). In the next phase, the contract pairs are generated from our seed datasets. Finally, we run three tools in pairs SoliDiffy , Difftastic , and Git line differencing tools according to our protocol for each research question on the generated contract pairs. The upcoming sections provide details of each stage of the experiment, including the dataset preparation, and the methodology for generating and processing source diff pairs.
IV-A Research Questions
The remaining sections of the paper address the following research questions:
  • RQ1: How does the performance of SoliDiffy compare to the most-closely related tool, Diffstastic?
  • RQ2: How does SoliDiffy perform when there are multiple changes in the smart contract source code?
  • RQ3: How does the type of syntactic changes in file affect the performance of SoliDiffy?
  • RQ4: How does SoliDiffy perform against line differencing on commit history of real-world smart contracts?
IV-B Protocol for Research Questions
IV-B1 Protocol for RQ1
To evaluate the effectiveness of SoliDiffy for AST differencing in Solidity code, we compare its performance against an existing open-source tool, Difftastic. We focus on the key metric of edit script length and success of the differencing task.
IV-B2 Protocol for RQ2
In this RQ, we analyze the effect of number of code differences on the performance of Solidity AST differencing tools, on the same dataset as RQ1.
IV-B3 Protocol for RQ3
The mutations used to generate the evaluation dataset range from simple syntax modifications to complex structural changes on Solidity smart contracts. We investigate the relationship between different operators and their edit distance.
IV-B4 Protocol for RQ4
We follow the same protocol as RQ1 with the difference of using a dataset of real-world commits in a popular smart contract project.
IV-C Datasets
IV-C1 Seed Datasets
To follow the protocols of RQ1–3, we need a dataset with a large number of Solidity source code files. For this, we use the DAppScan-source dataset [30, 32, 33]. This dataset consists of 39 9043990439\,90439 904 real-world Solidity Solidity source code files. From these 39 9043990439\,90439 904 files, $8102$ files are used for the our experiment in Figure 4, selected as follows: we remove all Solidity source files with duplicate names to ensure a more diverse dataset and reduce the potential for redundancy that could undermine the validity of our evaluation. The final dataset is available in a dedicated repository  [^2] https://github.com/SoliDiffy/SoliDiffyResults/tree/main/contracts/dataset.
For RQ4, we used commits of Uniswap v4 core smart contracts GitHub repository  [^3] https://github.com/Uniswap/v4-core.
IV-C2 Contract Pair Generation
To create contract pairs with varying levels of code alterations to the AST, one effective approach is to use a mutation testing tool [34]. Using code mutations for contract pair generation provides fine-grained code changes between versions, of better quality than what one may finds in repositories.
For this, we use mutation tool SuMo [35, 36]. It contains 44 mutation operators that are useful for evaluating various aspects of SoliDiffy [37, 38].
We used a script to invoke SuMo from the command line, generating mutants for all files in the dataset using each available mutation operator. The process involved iterating through all $44$ mutation operators, generating all possible mutations for each Solidity file, and creating up to $10$ mutated versions per file. In some cases, the actual number of generated mutants is lower than 10 due to the limited mutation opportunities in some files.
We note that some files in the dataset are incompatible with SuMo, causing crashes and preventing contract pair generation. In total, we generated 353 262353262353\,262353 262 contract pairs for differencing. The browsable version of these diff pairs is provided in our repository [^4] https://github.com/SoliDiffy/SoliDiffyResults/tree/master/contracts/mutants [^5] https://solidiffy.github.io/.
To generate contract pairs for differencing task of RQ4, we processed the commit history of Uniswap v4 core project. The contract pair generation begins by retrieving the entire commit history in chronological order using git log, followed by identifying the specific Solidity files altered in each commit through git diff-tree. For each modified file, the script extracts the version of the file at both the current and previous state and stores them. Then, a git diff between these two versions is computed using git diff, and the differences are saved to a file. This dataset of contract pairs and their differencing results are available publicly at our results repository  [^6] https://github.com/SoliDiffy/SoliDiffyResults/tree/master/uniswap-v4-diffs.
IV-D Execution Environment
We run the whole pipeline of the experiment on a system with an AMD EPYC 7742 64-core Processor and $528$ GB RAM. The total run time of the experiment was $6\mathrm{h}13\mathrm{m}43\mathrm{s}$. The differencing is parallelized based on the available number of CPU cores on the server.
V Experimental Results
V-A Results for RQ1
RQ1: How does the performance of SoliDiffy compare to the most-closely related tool, Diffstastic?
TABLE I: Effectiveness of Solidity differencing tools on our large dataset (Section IV-C1)
SoliDiffy Difftastic
Total diffed pairs 353 262353262353\,262353 262
Successful diffed pairs 339 596339596339\,596339 596 336 331336331336\,331336 331
[…]
We use SoliDiffy and Difftastic to conduct a large-scale campaign of Solidity smart contract source code differencing (see Section IV). Figure 5 shows the results of running the experiment, averaged across all diff pairs of each project (with varying contract pair modification severity and different types of modifications). We present the results as a side-by-side histogram. The green bars represent SoliDiffy, while the blue bars with a hatched pattern represent Difftastic. The y axis presents the frequency of edit script length that falls into a particular bin (x axis).
The key result is that SoliDiffy produces shorter edit scripts, as witnessed by the bars for SoliDiffy being consistently higher on the left side of the plot. Clearly, SoliDiffy produces fewer edit actions across most contracts of the dataset. In contrast, Difftastic’s distribution is more spread out, with some edit scripts containining more than $80$ changes and continuing to more than $500$ which were trimmed to fit the plot. To ensure that the visual observations are statisically significant, we performed a Wilcoxon signed-rank test [39] ($p<0.001$) that shows difference between edit script length pairs over all projects is statistically significant.
Moreover, as shown in Table I, SoliDiffy is able to successfully analyze more diff pairs than Diffstastic ($96.1\%$ vs. $95.1\%$). The main root cause of the crashes was syntax errors that were due to invalid syntax in mutated contracts.
Result for RQ1: SoliDiffy outperforms Difftastic by producing shorter edit scripts for Solidity smart contracts. Additionally, SoliDiffy successfully completed the analysis of a higher percentage of diffing tasks for contract pairs ($96.1\%$ vs. $95.1\%$).
V-B Results for RQ2
RQ2: How does SoliDiffy perform when there are multiple changes in the smart contract source code?
Figure 6: RQ2: Edit distances of SoliDiffy and Difftastic per initial number of mutations. Triangles annotate the average edit distance.
[…]
Figure 6 shows the results for the two differencing tools while running the diff pairs with the same operators repeatedly applied on the same Solidity contract (outlined in Section IV-C). The results are presented as a violin plot at which the width of the violin at different points shows the density of data. Key statistical markers, such as the mean and median in this plot highlight the skewness of the data. Peaks in the violin indicate where data clusters, and the tails represent outliers or extreme values. As Figure 6 shows, SoliDiffy is more dense towards the lower values of edit scripts, especially for lower number of mutations used for contract pair generation. For instance, differencing task performed on contract pairs which were the result of one and two mutations are very dense towards very low values of edit scripts for SoliDiffy.
Furthermore, consistent extreme peaks in Difftastic violins especially when having more number of mutations, shows it tends to generate very long edit scripts for at least a consistent proportion of differenced contract pairs especially when the number of modifications are increase.
While according to Figure 6 SoliDiffy exhibits lower values for the edit distance, we need a statistical test to confirm whether these differences are statistically significant. Given that the distributions represented in the violin plots are not normally distributed, we employed the Kolmogorov-Smirnov (K-S) test [40] to compare the two tools. The K-S test is suitable as it compares the entire cumulative distribution functions (CDFs) of two distributions, to detect differences not only in the central tendency but also in the overall shape, spread, and tails of the distributions. This is crucial because, as seen in the figure, Difftastic results show more variability and heavier tails compared to the more concentrated SoliDiffy distributions. The K-S test revealed statistically significant differences in all $10$ mutation severity comparisons, with p-values consistently below the conventional threshold of 0.05 with the highest p-value being $0.005$. The results of the our K-S test align with the visual representation in Figure 6, where SoliDiffy shows tighter distributions across all parameters, with the means consistently lower than Difftastic. SoliDiffy’s central tendency to lower edit distances, combined with the significantly different overall distribution shapes (as confirmed by the K-S test), provide compelling evidence of SoliDiffy’s superior performance.
Result for RQ2: SoliDiffy consistently produces smaller edit distances compared to Difftastic, regardless of the number of modifications in the smart contracts to be diffed. The finding holds statistical significance.
V-C Results for RQ3
RQ3: How does the type of syntactic changes in file affect the performance of SoliDiffy?
We analyze the results of applying diverse set of $44$ mutation operators on our dataset side-by-side for both SoliDiffy and Difftastic. Figure 7 presents the performance difference between the two tools. The exact mutation operator and its category are written bellow each bar pair. The bars show the effect of mutation operator and its category used to create the smart contract pairs. Each bar represents the mean edit distance between the original smart contract and its respective modified version produced by applying the specific mutation operator only once. For instance, when diff pairs consist of mutated code blocks category, that is, when large blocks of code are added, moved, or removed from the code, the Difftastic edit distances are significantly larger than SoliDiffy’s results or any other type of modification.
As Figure 7 presents, the performance of SoliDiffy and Difftastic is considerably different in many cases. In most cases that SoliDiffy performs better, it outperforms Difftastic by a great margin. For instance, in all differencing tasks belonging to the mutated code blocks category where full blocks of code are manipulated, SoliDiffy produces structurally meaningful edit scripts as opposed to Difftastic which tends to produce edit scripts consisting of word-by-word additions or deletions. SoliDiffy demonstrates a stronger performance where it matters most: in the cases where Difftastic falls short, the discrepancies are notably more pronounced, highlighting the superior efficiency of SoliDiffy in handling more complex differences, which are typical in real-world use-cases.
For the cases where SoliDiffy visibly produces longer edit scripts. For instance in the case of the textually minor mutation operator ICM (Increments Mirror) that changes an incrementing operator by swapping their two characters, += becomes =+. The problem when representing this change in AST edit actions is that the ASTs generated from these two versions are very different. In the first (+=), an operator is applied to two values and the new value is written to one of them. In the other (=+), a value is simply set to a negative value. Difftastic’s way of providing textual changes in this case provides a shorter edit distance as its edit script consists of adding + to an existing operand (=). Only the two characters that were swapped are displayed, and the resulting edit distance instead becomes the correct two per mutation. The same argument holds for all the instances that edit distances calculated by SoliDiffy is higher than Difftastic.
For diff pairs belonging to argument/modifier and miscellaneous categories, SoliDiffy and Difftastic inconsistently outperform each other. Our random sampling of diff pairs where Difftastic producing smaller edit distance confirms that these belong to the cases where Difftastic merely textually counts the number of add or removal of words in the Solidity contract.
We conducted a Wilcoxon signed-rank test [39] to assess the statistical significance of the observed differences between SoliDiffy and Difftastic across the $44$ mutation operators. The test results show that for all $44$ operators, the differences between SoliDiffy and Difftastic were statistically significantThis confirms that the visual performance differences seen in Figure 7 are not due to random chance.
Result for RQ3: SoliDiffy demonstrates superior performance in handling complex structural changes, particularly when large code blocks are modified. It can produce meaningful, concise edit scripts for diverse kinds of modification.
V-D Results for RQ4
RQ4: How does SoliDiffy perform against line differencing on commit history of real-world smart contracts?
Comparing Git line differencing which works on smart contract source level and not perform any structural (syntax-level) differencing, allows benchmarking SoliDiffy against the most simplistic way of transforming one smart contract to another by merely removing lines and adding new lines. We define the edit distance of Git line differencing results as the total number of lines marked to remove and add.
Figure 8 shows the results of differencing task on $925$ pairs of contract pairs with modifications from the commit history of Uniswap v4 core smart contracts. This figure visualizes SoliDiffy and Git line differencing results using a violin plot for each tool.
Figure 8: RQ4: Edit distances of SoliDiffy and Difftastic for $925$ smart contract pairs extracted from the commit history of Uniswap v4 core.
While the violin shapes indicate that the two tools perform similarly on the lower end of the distribution, the spread of values—particularly the tail behavior—differs slightly between the tools. SoliDiffy seems to consistently exhibit more compact results, while git diff demonstrates broader variability, especially in the higher ranges.
To check the statistical significance of observation in Figure 8, we conducted a Wilcoxon signed-rank test [39] on edit script length of two tools generated on the same contract pairs ($925$ pairs of edit distances). The result proved that distributions are indeed significantly different with a p-value of $0.01$ rejecting the null hypothesis that there is no statistically significant difference between these two edit script length distributions.
Result for RQ4: SoliDiffy works well to analyze open-source commits changing real-world Solidity smart contracts. SoliDiffy is readily usable by researchers and practitioners for advanced smart contract analysis tasks.
VI Discussion
VI-A Lessons Learned
Our experimental evaluation demonstrates that SoliDiffy offers significant improvements in edit script precision compared to Difftastic, particularly when handling complex code modifications, such as those involving large code blocks. By delivering shorter and more precise edit scripts, SoliDiffy provides developers with a clearer view of structural changes, reducing the cognitive load required for code reviews and audits. This is essential in the blockchain domain, where the immutability of deployed contracts necessitates rigorous pre-deployment analysis.
Our findings also revealed situations where SoliDiffy produced longer edit scripts than Difftastic. This was observed with certain low-impact syntax mutations where the distinction between AST nodes resulted in an increased edit distance. Such cases highlight that SoliDiffy’s more granular AST differencing may not always translate into shorter edit scripts, especially when the syntactic differences are minimal. Future work could explore hybrid approaches that incorporate both text-based and AST-based differencing to handle such cases more effectively.
The use of mutants as our primary dataset allowed for controlled evaluations, but it also introduced some limitations. Real-world smart contract updates often involve non-uniform changes that go beyond single syntactic mutations. The results of RQ4, using real-world data from Uniswap v4, suggest that SoliDiffy remains effective even in diverse commit histories, which indicates its robustness for practical applications. Further studies with varied real-world datasets could provide deeper insights into how well SoliDiffy performs in other scenarios, such as contract refactoring or collaborative development environments.
VI-B Threats to Validity
To evaluate SoliDiffy , we used mutants for generating diff pairs. This approach provided fine-grained control over the differences in the code, enabling us to systematically evaluate the tool (see Sections V-A, V-B, V-C). By using mutants instead of relying on the commit history of Solidity project repositories, we ensured that each diff pair was isolated and clearly attributable to specific types of syntactic changes. This level of control facilitated a more detailed analysis of the tools’ behavior under specific conditions.
However, using mutants also introduces some potential problems. Firstly, the differences between the resulting files may be identical across some casesFurthermore, many mutation operators generated minor changes, which, when analyzed individually, provided limited insight into the broader performance of the differencing tools. This could lead to an underestimation of the tools’ effectiveness in handling more complex, real-world code changes.
Even in cases where the large-scale evaluation showed differences in edit distances between the tools, these differences were not always apparent in the visual representation of the diffs.
Statistical analysis on edit distances also may introduce potential threats to validity. First, while non-parametric tests like the Wilcoxon signed-rank and Kolmogorov-Smirnov (K-S) tests are suitable for our non-normal data distributions, they assume comparable scales across tools. Differences in edit script lengths between SoliDiffy and Difftastic may partly reflect varying granularities rather than inherent tool efficacy. Second, the dataset’s synthetic nature (using controlled mutants) could bias statistical comparisons, as the tool behaviors observed in a laboratory setting may not fully extend to real-world contract complexities. Finally, the choice of statistical significance alone may overlook practical differences. While the tests confirm tool distinctions, additional metrics such as effect size or real-world user studies could provide further insights into the impact of these differences on developer tasks.
VII Related Work
TABLE II: Summary of notable tools and research on Solidity source code differencing.
Tool/Approach Key Features Differencing Technique Limitations Solidity
srcML [25] XML-based intermediate representation Unix diff on XML Limited evaluation and lacks structural analysis
Dex [41] Abstract Semantic Graphs (ASGs) instead of ASTs Graph differencing No move operation in edit scripts
UMLDiff [42] Reverse-engineered class models from Java code Structural analysis with similarity scores Limited to class structure, not applicable to all code changes
Diff/TS [43] Combines tree differencing with configurable heuristics Tree differencing No comparison to other similar tools in evaluation
OperV [20] Variable granularity using line and AST-based differencing Line and AST differencing Lacks comprehensive evaluation
MTDIFF [44] Optimizes edit script length Improved GumTree algorithm Comparable failure rates with other tools
IJM [45] Merges nodes, prunes sub-trees for faster differencing Improved GumTree algorithm Shorter, faster edit scripts
Matsumoto’s Approach [46] Splits AST nodes by line-based diff relevance Line and AST-based differencing Only focuses on improving specific troublesome actions
Yang & Whitehead’s Pruning [47] Pruning ASTs based on unchanged lines Line-based pruning for AST Pruning has negligible impact on diff results
HyperAST [24] Single AST representing multiple file versions AST storage across versions Limited to AST construction optimization
CLDiff [23] Groups and links related edit actions AST differencing with grouping More coarse-grained, focused on grouping related changes
SrcDiff [22] Heuristic-based matching, conversion rules Heuristic-based differencing Poor handling of complex updates in syntactic differencing
Difftastic [21] Supports Solidity Text changes differencing Lacks concrete evaluation, limited to text-based diffs, unsuitable for complex diff tasks, Supports only two edit actions
SoliDiffy (current work) Move operation in edit scripts, RTED algorithm, Solidity-specific differencing Top-down and bottom-up AST traversal and mapping to generate edit script for a diff pair, supports four edit actions Issues with visual representation of diffs, Edit script length for very small changes
As Solidity source code differencing is a contributing technical piece of many solutions such as version control [48], automated debugging [5], and code repair [49, 50], we provide a detailed review of tools and research on source code differencing.
VII-A Code Differencing
Tools such as srcML [25] and Dex [41] proposed ways to enhance code differencing, while UMLDiff [42] and Diff/TS [43] aimed at improving accuracy and efficiency. Meanwhile, OperV [20] explored varying levels of granularity in differencing.
The srcML tool converts source code into an XML-based intermediate representation, retaining both the syntax and textual elements, allowing for regenerating the original code [25]. Unlike AST-based tools, it uses the Unix diff command on these XML files.
Dex introduced the use of Abstract Semantic Graphs (ASGs) instead of ASTs for C code, linking related nodes like variable references and declarations [41]. Its differencing approach used graph rather than tree differencing, excluding the move operation from its edit scripts, but achieving a 95% accuracy in detecting correct edit actions.
UMLDiff [42] and Diff/TS [43] focused on enhancing the structural analysis of code changes. UMLDiff used class models reverse-engineered from Java source code to build change trees and calculated similarity scores to detect changes in the overarching class structure [42]. Diff/TS, on the other hand, combined tree differencing with configurable heuristics to improve runtime and accuracy, incorporating all standard edit actions in its scripts [43].
OperV sought to offer variable granularity in version control systems, combining line-based differencing with AST-based matching, though its evaluation was limited compared to more modern approaches [20].
The aforementioned tools do not provide the neither fine-grained differencing capabilities required for smart contract languages such as Solidity.
Recent AST differencing tools, including GumTree [19], MTDIFF [44], IJM [45], and the approach by Matsumoto et al. [46], focus on refining the generation of edit scripts (Section II-B2).
One common method for evaluating edit script quality is by measuring its length [19, 46, 44]. Shorter edit scripts are generally preferred because they tend to contain fewer redundant operations and more closely align with the actual code modifications. Another approach is to count the number of matched nodes in the initial differencing step, which are not included in the edit script, providing insight into the tool’s effectiveness in detecting unchanged code structures. However, there are criticisms of these methods. For instance, focusing solely on reducing script length can sometimes lead to suboptimal results, as seen in tools like SrcDiff [22].
In addition to quantitative measures, qualitative evaluations through expert analysis are also commonly used. These smaller-scale assessments, as applied in studies of tools like GumTree [19], Matsumoto’s approach [46], and the differential testing conducted by Fan et al. [51], provide insights into the real-world usefulness of AST differencing tools.
VII-B Gumtree Family of AST Differencing.
The Changedistiller algorithm [52] is a foundational work in AST differencing, introducing a method to match identical nodes between two ASTs and generate an edit script. It built upon Chawathe’s 1996 algorithm [28], optimizing it for source code by reducing edit script length by 45%. This approach influenced many subsequent tools, including GumTree [19].
GumTree is particularly notable for its introduction of the move operation in edit scripts, which improves accuracy by grouping related changes. It uses a combination of top-down and bottom-up AST traversal and incorporates the RTED algorithm [53] for generating mappings in smaller sub-trees. However, GumTree has not integrated the more recent APTED algorithm [54], which could enhance its performance.
GumTree also supports hyperparameter tuning to optimize edit script length, as demonstrated by Martinez’s Diff Auto Tuning (DAT) technique, which reduced script length [55]. Additionally, GumTree can process general-purpose Tree-sitter CSTs by converting them into a format suitable for AST differencing [56].
MTDIFF [44] and Iterative Java Matcher (IJM) [45] introduced improvements on built on top of core GumTree algorithm but lack maturity and integration into the mainstream differencing code tool.
While the aforementioned tools offer improvements for general-purpose languages, they lack specific adaptations for Solidity. SoliDiffy builds on this line of work by building on top of Gumtree’s algorithms and ecosystem for Solidity smart contracts.
VII-C Solidity Code Differencing
Research on code differencing specific to Solidity is limited. While line-based differencing can be used across languages, it lacks the precision needed for Solidity’s unique syntax and semantics, as discussed in Section II-B.
The only dedicated Solidity differencing tool in the literature is part of the Solidity Instrumentation Framework (SIF) [57]. SIF uses AST-differencing, but its implementation is poorly documented and relies on an outdated AST format no longer supported by the Solidity compiler, making it unusable for newer code.
Outside academic literature, Smartdiffer [58] links Etherscan with a line-based differencing tool but does not perform AST differencing, limiting its relevance to more sophisticated analysis.
Difftastic [21] is a more complete AST differencing tool that supports Solidity. It uses a tree-sitter parser to generate side-by-side diffs or JSON output. However, it does not use traditional edit script generation, instead focusing on concrete text changes, and lacks thorough research-based evaluation.
VIII Conclusion
We introduced SoliDiffy, a novel AST differencing tool for Solidity smart contracts. SoliDiffy provides fine-grained, accurate differencing, outperforming existing tools in both edit script quality and ability to handle complex syntactic changes. Our evaluation demonstrated that SoliDiffy supports AST differencing of complex changes and real world contracts. SoliDiffy also gives an intuitive diff representation for developers. SoliDiffy sets a solid foundation for future enhancements in the field of smart contract analysis, such as incorporating semantic analysis or extending support to other blockchain languages.